HUBBLE SPACE TELESCOPE DAILY REPORT #3106 PERIOD COVERED: 0000Z (UTC) 04/30/02 - 0000Z (UTC) 05/01/02 Daily Status Report as of 121/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Five Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.2 Completed Two Sets of STIS/CCD 8902 (Dark Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. No problems were reported. 1.3 Completed STIS/MA2 9573 (NUV-MAMA Daily Dark Monitor) Abstract The Space Telescope Imaging Spectrograph (CCD and MA2) was used to perform daily monitoring of the NUV MAMA detector dark noise, monitoring the effects of thermal changes on the NUV dark rate. No anomalous activity was noted. 1.4 Completed Two Sets of NICMOS/1/2/3 8944 (Filter Wheel/Mechanisms Mini-Functional Test) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3) was used to perform an early engineering test to verify the aliveness, functionality, operability, and electro-mechanical calibration of the NICMOS filter wheel motors and assembly. This was the first use of the NICMOS filter wheel mechanisms since they were disabled by ground command in January, 1999. This test was designed to obviate concerns over possible deformation or breakage of the fitter wheel "soda-straw" shafts due to excess rotational drag torque and/or bending moments which may be imparted due to changes in the dewar metrology from warm-up/cool-down. No anomalous incident occurred. 1.5 Completed STIS/MA1 8922 (FUV-MAMA Cycle 10 Flats) The Space Telescope Imaging Spectrograph (MA1) was used to obtain FUV-MAMA observations of the STIS internal Krypton lamp to construct an FUV flat applicable to all FUV modes. No problems were encountered. 1.6 Completed Three Sets of WF/PC-2 8937 (Cycle 9 Supplemental Darks pt2/3) The WF/PC-2 was used obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. No problems were encountered. 1.7 Completed Two Sets of ACS/WFC 9029 (Grism/Prism Performance Check) The Advance Camera for Surveys (WFC) was used to observe a Wolf-Rayet star and a White Dwarf the grism of both the Wide Field and eventually the High Resolution Channels to measure: 1} the dispersion of the grism and its field dependence; 2} the grism throughput and its field dependence; 3} the frequency of the flat field variation {L-flat} as a function of wavelength; 4} the fringe pattern at longer wavelengths as a function of position on the chip. These measurements will be carried out at several positions on the chip including the centre and the corners of the WFC and HRC chips. All observations completed without incident. 1.8 Completed STIS/CCD 8904 (Bias Monitor-Part 2) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 in order to build up high-S/N superbiases and track the evolution of hot columns. The proposal completed with no anomalous activity. 1.9 Completed Four Sets of WF/PC-2 9318 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform the generic target version of the WFPC2 Archival Pure Parallel program. The program was used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. There were no reported problems. 1.10 Completed ACS/WFC 9587 (HST Imaging Polarimetry of the Light Echo around V838 Monocerotis) The Advanced Camera for Surveys (WFC) was used to observe V838 Mon0cerotis which is a completely unanticipated new kind of object: although it has undergone an outburst somewhat similar to that of an extremely slow nova, its spectrum is unique and totally unlike that of any type of nova. Knowledge of its distance and luminosity is crucial in understanding its nature and origin. Remarkably, a rapidly evolving light echo around V838 Monocerotis has been detected from the ground in recent weeks. HST polarimetric imaging over the next few months, as the light echo expands and fades, will provide a direct geometrical distance to this object. Since the only previous Galactic nova light echoes occurred in 1901 and 1936, this extraordinary combination of circumstances is unlikely to recur during the HST mission. All observations completed with no reported problems. 1.11 Completed Two Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3) was used 1) to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 1.12 Completed STIS/MA1 8920 (Cycle 10 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to perform the routine monitoring of the MAMA detector dark noise, and is the primary means of checking on health of the MAMA detectors systems through frequent monitoring of the background count rate. The proposal completed with no reported anomalous activity. 1.13 Completed STIS/CCD 9050 (Outflow Collimation in Bipolar Symbiotic Nebulae) The Space Telescope Imaging Spectrograph (CCD) was used to observe flow collimation in evolved stars that is neither expected nor understood. Classical theories of stellar evolution do not predict and cannot explain this bipolarity. More exotic concepts {binary interactions, spun-up atmospheres, poloidal or toroidal magnetized winds} have been proposed, but observations are yet to verify or falsify any of their predictions. This proposal will probe the near-nuclear morphology and kinematics of four bright, low-extinction targets whose large-scale structure is highly bipolar. The goal is to provide a detailed description of the circumnuclear outflows, to uncover the physical structure and nature of the collimator, and to evaluate the speculative collimation mechanisms. The bright nucleus has hampered efforts to explore the nebular collimators that lie close to the star, so we'll use STIS to disperse the nuclear light and, thus, to avoid its glare. A secondary goal is to obtain second-epoch WFPC2 images of all targets. There were no reported problems. 1.14 Completed Two Sets of STIS/CCD 9088 (Next Generation Spectral Library of Stars) The Space Telescope Imaging Spectrograph (CCD) was used to produce a "Next Generation'' Spectral Library of 600 stars for use in modeling the integrated light of galaxies and clusters by using the low dispersion UV and optical gratings of STIS. The library will be roughly equally divided among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling the entire HR-diagram in each bin. Such a library will surpass all extant compilations and have lasting archival value, well into the Next Generation Space Telescope era. No problems occurred. 1.15 Completed Three Sets of STIS/CCD 9317 (Pure Parallel Imaging Program: Cycle 10) The Space Telescope Imaging Spectrograph (CCD) was used to perform the default archival pure parallel program for STIS during cycle 10. There were no reported problems. 1.16 Completed ACS/WFC 9575 (Default {Archival} Pure Parallel Program) The Advanced Camera for Surveys (WFC) was used to test ACS pure parallels in POMS. There were no reported problems. 1.17 Completed NIC/1/2/3 8974 (Flats and Temperature Dependence of the DQE) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3) was used to obtain initial estimates of the detective quantum efficiency {DQE} of the NICMOS detectors and its temperature dependence in the previously uncharted temperature regime expected for operation under the NICMOS Cooling System {NCS}. The observations measured the relative and absolute DQE variation at three temperature setpoints. In addition, they provided a monitor for particulate contamination and detector lateral position. When stars are present in the field of view, they will enable a preliminary focus determination. No problems were encountered. 1.18 Completed STIS/CCD 9077 (Survey of the LMC Planetary Nebulae) The Space Telescope Imaging Spectrograph (CCD) was used to perform a snapshot survey of all known LMC planetary nebulae {PNe} in order to study the co-evolution of the nebulae and their central stars, and to probe the chemical enrichment history of the LMC. There were no reported problems. 1.19 Completed STIS/CCD/MA2 9051 (Identifying Damped Lyman-alpha Galaxies at z~1) The Space Telescope Imaging Spectrograph (CCD and MA2) was used to look for damped Lyman-alpha absorption systems that contain the bulk of the neutral gas in the Universe in the redshift range z = 0.5 - 5, yet the nature of the galaxies responsible for the absorption is not well understood. Only recently have observers found more than a handful of damped absorbers at redshifts z < 1.5. Using the FIRST Bright Quasar Survey {FBQS}, with over a 1000 quasars, the proposers have undertaken a survey to build a complete picture of he nature of the galaxies responsible for damped Lyman-alpha absorption systems at z~1 and to double the sample size at this redshift. No problems were encountered. 1.20 Completed WF/PC-2 9319 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to execute a POMS test proposal, designed to simulate future scientific plans. The proposal completed without incident. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 13 Successful: 13 Scheduled Re-acquisitions: 2 Successful: 2 2.2 FHST Updates: Scheduled: 27 Successful: 27 2.3 Operations Notes: The SSR-3 EDAC error counter was cleared at 120/1422Z and at 120/1810Z per ROP SR-9A. Using an operations request, the NICMOS Cooling System alternate PID Setpoint was adjusted by 0.5 degK to 72.5 degK at the weighted average neon temperature at 120/1840Z. The compressor speed ramped back to approximately 6890 rps to accommodate the transition to the new control point temperature. Another NICMOS filter wheel test was concluded at 120/2238Z. Analysis of the data indicates that all functions performed as expected with results closely approximating those obtained in 1997. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW